OpenStudy (boxman61):

Rolls of aluminum foil are 305 mm wide and 0.015 mm thick. What maximum length of aluminum foil can be made from 0.83 kg of aluminum?( d(aluminum)=2.70 g/cm3)

OpenStudy (evoker):

Turn the mass into a volume using the density and then use the volume equation to find the length.

OpenStudy (evoker):

Though you will need to convert everything so that the units match.

OpenStudy (boxman61):

V=M/D D=2.70g/cm^3 M=0.83kg \[\left(\begin{matrix}0.83Kg \\ 1\end{matrix}\right)*\left(\begin{matrix}1000g \\ 1kg\end{matrix}\right)=830g\] \[\left(\begin{matrix}830g \\ 270g/cm^3\end{matrix}\right)=307.41cm^3\]

OpenStudy (evoker):

Looks good so far.

OpenStudy (boxman61):

im stuck here...how do i account for width and thickness?

OpenStudy (evoker):

V=LWH

OpenStudy (evoker):

Where H would be the thickness

OpenStudy (boxman61):

to confirm \[307.41=L*305mm*0.015mm\]

OpenStudy (evoker):

Not quite once again you will need to do conversions

OpenStudy (evoker):

your volume is cm^3 while those measurements are in mm

OpenStudy (boxman61):

\[305mm \rightarrow 305000mm^3 \rightarrow 305cm^3\] \[0.015mm \rightarrow 15mm^3 \rightarrow 0.015cm^3\] \[307.41cm^3=L*305cm^3*0.015cm^3\] L=67.2cm^3

OpenStudy (boxman61):

67.2cm^3 = 4.07 cm = 0.0407meters

OpenStudy (evoker):

Conversions appear to be off, we can put it all in cm and cm^3 if you like

OpenStudy (evoker):

to go from mm to cm you would divide by 10.

OpenStudy (evoker):

and then you can leave the volume alone

OpenStudy (boxman61):

so 305mm /10 =30.5cm and 0.015mm /10 = 0.0015cm

OpenStudy (evoker):

Yes

OpenStudy (boxman61):

307.41cm^3 = 6.75cm 6.75=L*30.5*0.0015 L=147.54cm or 1.48 Meters

OpenStudy (evoker):

you can leave the volume as it is, it is already in cm^3.

OpenStudy (evoker):

307.41=L*30.5*.0015 where L will be in cm

OpenStudy (boxman61):

answer needs to be in meters so wouldn't i need to convert the cm^3 to cm?

OpenStudy (evoker):

well when you are finished I would convert the length to meters by dividing by 100.

OpenStudy (evoker):

You could turn everything to meters but much easier to just solve for L and then convert that answer to meters.

OpenStudy (boxman61):

Oh I see because L W and H are all cm when they multiply the become cm^3 so If I do 307.41=L*30.5*.0015 L=6719.34

OpenStudy (evoker):

Yes

OpenStudy (boxman61):

67.2 meters

OpenStudy (evoker):

Looks good

OpenStudy (boxman61):

Whew... thank you very much for sticking with me on that one. I really appreciate it

OpenStudy (evoker):

No problem, all the conversions do make it a bit tricky